OpenStudy (anonymous):

Given the equation y − 4 = (x + 8) in point-slope form, identify the equation of the same line in standard form

OpenStudy (nubeer):

a standard equation is in form of y = mx +b so solve the given equation.. y - 4 = x +8 add +4 on both sides it will standardised ur equation

OpenStudy (cwrw238):

hhmm isn't that slope intercept form ? (y = mx + b)

OpenStudy (anonymous):

yeaa

OpenStudy (anonymous):

ohh